# Comprehending Knowledge Graphs with Large Language Models for Recommender Systems

### v2

@@ Abstract.
In recent years, the introduction of knowledge graphs (KGs) has significantly advanced recommender systems by facilitating the discovery of potential associations between items. However, existing methods still face several limitations. First, most KGs suffer from missing facts or limited scopes. Second, existing methods convert textual information in KGs into IDs, resulting in the loss of natural semantic connections between different items. Third, existing methods struggle to capture high-order connections in the global KG. To address these limitations, we propose a novel method called CoLaKG, which leverages large language models (LLMs) to improve KG-based recommendations. The extensive world knowledge and remarkable reasoning capabilities of LLMs enable our method to supplement missing facts in KGs. Additionally, their powerful text understanding abilities allow for better utilization of semantic information. Specifically, CoLaKG extracts useful information from the KG at both local and global levels. By employing item-centered subgraph extraction and prompt engineering, it accurately captures the local KG. Subsequently, through retrieval-based neighbor enhancement, it supplements the current item by capturing related items from the entire KG, thereby effectively utilizing global information. The local and global information extracted by the LLM are effectively integrated into the recommendation model through a representation fusion module and a retrieval-augmented representation learning module, respectively, thereby improving recommendation performance. Extensive experiments on four real-world datasets demonstrate the superiority of our method.
Knowledge Graphs, Large Language Models, Recommendation
ccs: Information systems Recommender systems
@@ 1. Introduction
The rapid advancement of online platforms has led to an increasingly critical issue of information overload. Recommender systems address this problem by modeling user preferences based on historical data. Collaborative filtering (CF) (Rendle et al., 2012; He and Chua, 2017; He et al., 2020), as one of the most classic and efficient methods, has been extensively employed in existing recommender systems. However, CF-based methods exclusively rely on user-item interaction records, often suffering from the data sparsity issue (Togashi et al., 2021; Cui et al., 2024). To address this issue, recent studies (Wang et al., 2019a; Zhang et al., 2016; Yang et al., 2022) have incorporated knowledge graphs (KGs) as external knowledge sources into recommendation models, achieving significant progress. Typically, these methods capture diverse and higher-order relationships between items by modeling the structure and attribute information in KGs, thereby enhancing the learning process of user and item representations (Yang et al., 2022).
@@
Despite the effectiveness of existing KG-enhanced recommendation methods, they still have several limitations. i) First, many KGs suffer from missing facts and limited scopes (Guo et al., 2020), as constructing KGs often requires manual effort and domain expertise. The absence of key attributes, such as the genres of a movie, can cause items that originally share the same attribute to lose their semantic connections. As illustrated in Figure 1, items A and B should have a connecting path (A-P-B) in the KG. However, due to item A missing the P attribute, A and B are now not associated with each other. Such biased knowledge will lead to suboptimal performance of the recommendation model. ii) Second, existing methods (Yang et al., 2022, 2023; Wang et al., 2019a, 2021, 2024) convert textual entities and relations into IDs for use, failing to leverage semantic information inherent in the text. Moreover, this can result in losing natural semantic connections between items. For instance, in Figure 1, “horror” and “thriller” are two semantically related attributes of Item F and Item G, respectively. However, similar semantics are not reflected in different entity IDs (51 and 320), which further results in the originally related items F and G having no connections in the KG. iii) Third, existing methods (Wang et al., 2019a; Yang et al., 2022, 2023; Wang et al., 2021, 2024; Tian et al., 2021) struggle to capture high-order relationships in global KGs. Most of them propagate and aggregate information by stacking multiple layers of graph neural networks (GNNs). The layer-by-layer propagation is not only inefficient but also accumulates a large amount of irrelevant node information, leading to the over-smoothing issue (Guo et al., 2020; Wang et al., 2024). For instance, assumming that points A and H in Figure 1 have a strong semantic connection, the considerable distance between them in the KG presents big challenges for existing methods in capturing this semantic relation. Due to the aforementioned limitations, existing KG-based recommendation methods are unable to identify the connections between historically interacted items and target items in the knowledge graph, resulting in suboptimal predictions. This is illustrated in the lower left corner of Figure 1.
@@
Empowered by extensive knowledge and remarkable reasoning abilities, large language models (LLMs) have demonstrated significant promise in semantic understanding and knowledge extraction. Consequently, LLMs have the potential to address the aforementioned issues. Recently, many studies have leveraged LLMs to improve recommendation models. For example, some studies utilize LLMs to generate semantic representations of item profiles (Ren et al., 2024). Other studies (Zhao et al., 2024) employ LLMs to determine whether a complementary relationship exists between two items, thereby recommending complementary products based on users’ historical behaviors. However, these methods do not fully exploit the semantic and structural information of KGs. As one of the most common and important sources of knowledge, KGs contain a wealth of semantic associations among entities and relations, which are often overlooked by existing methods that typically consider only item profiles. Additionally, KGs serve as task-relevant knowledge repositories, effectively aiding LLMs in acquiring task-specific knowledge and mitigating the issue of hallucinations caused by excessive divergence. Nevertheless, effectively leveraging LLMs to handle the diverse semantic relationships in KGs is highly challenging, as it is unrealistic to input the vast number of entities and relationships in KGs into an LLM.
To bridge this gap, in this paper, we propose a novel method named Comprehending Knowledge Graphs with Large Language Models for Recommendation (CoLaKG). The core idea is to leverage LLMs for understanding the semantic and structural information of KGs to enhance the representation learning of users and items. Our method comprises two stages:
  • Comprehending KGs with LLMs. Given the impracticality of inputting the entire KG into LLMs, we propose leveraging KG information by two components: 1) Local KG Comprehension, which involves extracting item-centered KG subgraphs for analysis by LLMs, and 2) Retrieval-based Global KG Utilization, which utilizes the semantic similarity of KG subgraphs to retrieve semantically related items from the entire KG, thereby leveraging global KG information.
@@
  • Incorporating semantic embeddings into the recommendation model. Our objective is to integrate the semantic embeddings of KGs with the ID embeddings of the recommendation model, thereby leveraging both collaborative signals and semantic information. First, we align and fuse the item ID representations with the item-centered KG subgraph representations. Then, we employ a retrieval-augmented representation learning method to further enhance item representations with semantically similar neighbors from the global KG, thereby incorporating valuable information from the global KG.
@@
It is important to note that these two stages are decoupled, meaning that our model does not involve LLM inference during the recommendation process. This decoupling allows our model to be effectively applied in real-world recommendation scenarios.
@@ 2.1. Knowledge-aware Recommendation
Existing knowledge-aware recommendation methods can be categorized into three types (Guo et al., 2020): embedding-based methods, path-based methods, and GNN-based methods. Embedding-based methods (Cao et al., 2019; Zhang et al., 2016; Wang et al., 2018b) enhance the representations of users and items by leveraging the relations and entities within the KGs. Notable examples include CKE (Zhang et al., 2016), which integrates various types of side information into a collaborative filtering framework using TransR (Lin et al., 2015) for structural knowledge. Another example is DKN (Wang et al., 2018b), which improves news representations by combining textual embeddings of sentences and knowledge-level embeddings of entities. Path-based methods leverage KGs to explore long-range connectivity (Wang et al., 2019b; Yu et al., 2014; Hu et al., 2018). For example, Personalized Entity Recommendation (PER) (Yu et al., 2014) treats a Knowledge Graph (KG) as a heterogeneous information network and extracts meta-path-based latent features to represent the connectivity between users and items along various types of relational paths. MCRec (Hu et al., 2018) constructs meta-paths and learns the explicit representations of meta-paths to depict the interaction context of user-item pairs. Despite their effectiveness, these approaches heavily rely on domain knowledge and human effort for meta-path design. Recently, GNN-based methods have been proposed, which enhance entity and relation representations by aggregating embeddings from multi-hop neighbors (Wang et al., 2019c, a, 2021; Liang et al., 2023; Chen et al., 2022). For instance, KGAT (Wang et al., 2019a) employs graph attention mechanisms to propagate embeddings and utilizes multi-layer perceptrons to generate final recommendation scores in an end-to-end manner. Similarly, KGIN (Wang et al., 2021) adopts an adaptive aggregation method to capture fine-grained user intentions. Additionally, some methods (Yang et al., 2022, 2023; Wang et al., 2024; Zhu et al., 2023) employ contrastive learning to mitigate potential knowledge noise and identify informative knowledge connections.
@@ 2.2. LLMs for Recommendation
In light of the emergence of large language models and their remarkable achievements in the field of NLP, scholars have begun to explore the potential application of LLMs in recommender systems (Wu et al., 2023; Zhao et al., 2023b; Fan et al., 2023; Chen et al., 2023). Due to the powerful reasoning capabilities and extensive world knowledge of LLMs, they have been already naturally applied to zero-shot (Hou et al., 2024; He et al., 2023; Wang and Lim, 2023) and few-shot recommendation scenarios (Bao et al., 2023; Lin et al., 2024). In these studies, LLMs are directly used as a recommendation model (Zhao et al., 2023a; Li et al., 2023), where the output of LLMs is expected to offer a reasonable recommendation result (Wu et al., 2024b). However, when the dataset is sufficiently large, their performance often falls short of that achieved by traditional recommendation models. Another line of research involves leveraging LLMs as feature extractors. These methods (Ren et al., 2024; Zhu et al., 2024; Hu et al., 2024; Wu et al., 2024a; Acharya et al., 2023; Kim et al., 2024; Wei et al., 2024; Ren et al., 2024) generate intermediate decision results or semantic embeddings of users and items, which are then input into traditional recommendation models to produce the final recommendations. Unlike existing methods, our approach aims to leverage the extensive knowledge base and reasoning capabilities of LLMs to understand KGs and transform them into semantic embeddings, thereby addressing existing issues in KG-based recommender systems and enhancing recommendation performance.
@@ 3. Preliminaries
  • Knowledge Graph. We capture real-world knowledge about items using a heterogeneous graph composed of triplets, represented as $\mathcal{G}_{k}=\{(h,r,t)\}$. In this context, $h$ and $t$ are knowledge entities belonging to the set $\mathcal{E}$, while $r$ is a semantic relation from the set $\mathcal{R}$ that links them, as exemplified by the triplet (James Cameron, directed, Titanic). Notably, the item set is a subset of the entity set, denoted as $\mathcal{V}\subset\mathcal{E}$. This form of knowledge graph enables us to model the intricate relationships between items and entities.
@@ 4.1.2. Rerieval-based Global KG Utilization
This section introduces the utilization of global KG information. In recommendation scenarios, items that are distant in the KG can still have close semantic associations. However, the number of KG nodes increases exponentially with the number of hops, and the input length of LLMs is limited by the number of tokens. This makes it impractical to input the entire KG into an LLM directly. To address this challenge, we propose a simple and effective method called retrieval-based global KG utilization.
@@
From a high-level perspective, we transform higher-order associations between items in the KG into semantic relationships on the constructed item-item graph $\mathcal{G}_{v}$. Based on this foundation, it is essential to retrieve items that are semantically strongly related to the given item, as items with lower semantic relevance may introduce noise. Specifically, given an item $v_{i}$, we rank all other items $v_{j}\in\mathcal{V}$ where $v_{j}\neq v_{i}$ in descending order based on the semantic similarity $r_{(v_{i},v_{j})}$. Subsequently, we retrieve the top-$k$ items with the highest similarity scores, forming the neighbor set of $v_{i}$: $\mathcal{N}_{k}(v_{i})$, where $0<k<|\mathcal{V}|$ is an adjustable hyperparameter, representing the number of retrieved neighbors.
In this manner, we filter out items with low semantic associations to the current item in the entire KG. Traditional KG-based recommendation methods aggregate related items through layer-by-layer information propagation on the KG. High-order item associations require many propagation and aggregation steps to be captured. In contrast, our retrieval-based method can recall strongly semantically related neighbors of any order across the entire graph through a single semantic similarity calculation. The purpose of retrieving semantically related neighbors is to leverage them to enhance item representations, which will be introduced in Section 4.3.2.
@@ 4.2. User Preference Comprehension
4.3. Retrieval-Augmented Representation Learning
@@ 4.4. User-Item Modeling
where $\hat{y}_{uv}$ is the predicted probability of user $u$ interacting with item $v$, $\mathbf{h}_{u}$ is the representation for user $u$, $\mathbf{h}_{v}$ is the augmented representation for item $v$, and $\mathcal{F}$ denotes the function of the recommendation model.
@@ 4.5. Model Training
Our approach can be divided into two stages. In the first stage, we employ the LLM to comprehend the KGs, generating corresponding semantic embeddings for each item and user, denoted as $\mathbf{s}_{v}$ and $\mathbf{s}_{u}$, respectively. In the second stage, these semantic embeddings are integrated into the recommendation model through an adapter network to enhance its performance. Only the second stage necessitates supervised training, where we adopt the widely-used Bayesian Personalized Ranking (BPR) loss:

### v3

@@ Abstract.
In recent years, the introduction of knowledge graphs (KGs) has significantly advanced recommender systems by facilitating the discovery of potential associations between items. However, existing methods still face several limitations. First, most KGs suffer from missing facts or limited scopes. Second, existing methods convert textual information in KGs into IDs, resulting in the loss of natural semantic connections between different items. Third, existing methods struggle to capture high-order connections in the global KG. To address these limitations, we propose a novel method called CoLaKG, which leverages large language models (LLMs) to improve KG-based recommendations. The extensive knowledge and remarkable reasoning capabilities of LLMs enable our method to supplement missing facts in KGs, and their powerful text understanding abilities allow for better utilization of semantic information. Specifically, CoLaKG extracts useful information from KGs at both local and global levels. By employing the item-centered subgraph extraction and prompt engineering, it can accurately understand the local information. In addition, through the semantic-based retrieval module, each item is enriched by related items from the entire knowledge graph, effectively harnessing global information. Furthermore, the local and global information are effectively integrated into the recommendation model through a representation fusion module and a retrieval-augmented representation learning module, respectively. Extensive experiments on four real-world datasets demonstrate the superiority of our method. The code of our method is available at https://github.com/ziqiangcui/CoLaKG.
Large Language Models, Knowledge Graphs, Recommendation
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 13–18, 2025; Padua, Italybooktitle: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR ’25), July 13–18, 2025, Padua, Italydoi: 10.1145/3726302.3729932isbn: 979-8-4007-1592-1/2025/07ccs: Information systems Recommender systems
@@ 1. Introduction
The rapid advancement of online platforms has led to an increasingly critical issue of information overload. Recommender systems address this problem by modeling user preferences based on historical data. Collaborative filtering (CF) (Rendle et al., 2012; He and Chua, 2017; He et al., 2020), as one of the most classic and efficient methods, has been extensively employed in existing recommender systems. However, CF-based methods exclusively rely on user-item interaction records, often suffering from the data sparsity issue (Togashi et al., 2021). To address this issue, recent studies (Wang et al., 2019a; Zhang et al., 2016; Yang et al., 2022) have incorporated knowledge graphs (KGs) as external knowledge sources into recommendation models, achieving significant progress. Typically, these methods capture diverse and high-order relationships between items by modeling the structure and attribute information in KGs, thereby enhancing the learning process of user and item representations (Yang et al., 2022).
@@
Despite the effectiveness of existing KG-enhanced recommendation methods, they still face several challenges. i) First, many KGs suffer from missing facts and limited scopes (Guo et al., 2020; Wang et al., 2023), as constructing KGs often requires manual effort and domain expertise. The absence of key attributes, such as the genres of a movie, can cause items that originally share the same attribute to lose their connections. As illustrated in Figure 1, items A and B should have a connecting path (A-P-B) in the KG. However, due to item A missing the P attribute, A and B are now not associated with each other. Such biased knowledge will lead to suboptimal performance of the recommendation model. ii) Second, existing methods (Yang et al., 2022, 2023; Wang et al., 2019a, 2021, 2024) convert textual entities and relations into IDs for use, which can result in losing natural semantic connections between items. For instance, in Figure 1, “horror” and “thriller” are two semantically related attributes of Item F and Item G, respectively. However, similar semantics are not reflected in different entity IDs (51 and 320), which further results in the related items F and G having no connections in the KG. iii) Third, existing methods (Wang et al., 2019a; Yang et al., 2022, 2023; Wang et al., 2021, 2024; Tian et al., 2021) struggle to capture high-order relationships in the entire KG. Most of them propagate and aggregate information by stacking multiple layers of graph neural networks (GNNs). The layer-by-layer propagation is not only inefficient but also accumulates a large amount of irrelevant node information, leading to the over-smoothing issue (Guo et al., 2020; Wang et al., 2024). For instance, assumming that points A and H in Figure 1 have a strong semantic connection, the considerable distance between them in the KG presents big challenges for existing methods in capturing this relation. As shown in the lower left corner of Figure 1, due to the aforementioned three limitations, existing methods fail to identify the potential connections between historically interacted items and the target item through the KG, resulting in suboptimal recommendations.
@@
Empowered by extensive knowledge and remarkable reasoning abilities, large language models (LLMs) have demonstrated significant promise in semantic understanding and knowledge extraction. Recently, many studies have leveraged LLMs to improve recommendation models. For example, some studies utilize LLMs to generate semantic representations of item profiles (Ren et al., 2024), while other studies (Zhao et al., 2024) employ LLMs to determine whether a complementary relationship exists between two items, thereby recommending complementary products based on users’ historical behaviors. However, these methods do not fully exploit the semantic and structural information of KGs. As one of the most common and important sources of knowledge, KGs contain a wealth of semantic associations among entities and relations, which are often overlooked by existing methods that typically consider only item profiles. Additionally, KGs serve as task-relevant knowledge repositories, effectively aiding LLMs in acquiring task-specific knowledge and mitigating the issue of hallucinations caused by excessive divergence. Nevertheless, leveraging LLMs to handle the diverse semantic relationships in KGs is highly challenging, as it is unrealistic to input the vast number of entities and connections in KGs into a language model.
@@
To bridge this gap, in this paper, we propose a novel method named Comprehending Knowledge Graphs with Large Language Models for Recommendation (CoLaKG). The core idea is to leverage LLMs for understanding the semantic and structural information of KGs to enhance the representation learning of items and users. Our method comprises two stages: i) Comprehending KGs with LLMs. Given the impracticality of inputting the entire KG into an LLM, CoLaKG leverages KG information by two components: a Local KG Comprehension module which involves extracting item-centered KG subgraphs for analysis by LLMs, and a Retrieval-based Global KG Utilization module which retrieves semantically related items from the entire KG to effectively leverage global KG information. ii) Incorporating semantic embeddings into the recommendation model. This stage integrates the semantic embeddings of KGs with the ID embeddings, thereby leveraging both collaborative signals and semantic information. First, we align and fuse the item ID representations with the item-centered KG subgraph representations. Then, we employ a retrieval-augmented representation method to further enhance item representations with semantically similar neighbors from the entire KG.
@@
It is important to note that these two stages are decoupled, meaning that our method does not involve LLM inference during the recommendation process. This allows our method to be efficiently applied in real-world recommendation scenarios.
@@ 2.1. Knowledge-aware Recommendation
Existing knowledge-aware recommendation methods can be categorized into three types (Guo et al., 2020): embedding-based methods, path-based methods, and GNN-based methods. Embedding-based methods (Cao et al., 2019; Zhang et al., 2016; Wang et al., 2018b; Lin et al., 2024a) enhance the representations of users and items by leveraging the relations and entities within the KGs. Notable examples include CKE (Zhang et al., 2016), which integrates various types of side information into a collaborative filtering framework using TransR (Lin et al., 2015). Another example is DKN (Wang et al., 2018b), which improves news representations by combining textual embeddings of sentences and knowledge-level embeddings of entities. Path-based methods leverage KGs to explore long-range connectivity (Wang et al., 2019b; Yu et al., 2014; Zhang et al., 2024; Hu et al., 2018). For example, Personalized Entity Recommendation (PER) (Yu et al., 2014) treats a Knowledge Graph (KG) as a heterogeneous information network and extracts meta-path-based latent features to represent the connectivity between users and items along various types of relational paths. MCRec (Hu et al., 2018) constructs meta-paths and learns the explicit representations of meta-paths to depict the interaction context of user-item pairs. Despite their effectiveness, these approaches heavily rely on domain knowledge and human effort for meta-path design. Recently, GNN-based methods have been proposed, which enhance entity and relation representations by aggregating embeddings from multi-hop neighbors (Wang et al., 2019c, a, 2021). For instance, KGAT (Wang et al., 2019a) employs the graph attention mechanism to propagate embeddings and utilizes multi-layer perceptrons to generate final recommendation scores in an end-to-end manner. Similarly, KGIN (Wang et al., 2021) adopts an adaptive aggregation method to capture fine-grained user intentions. Additionally, some methods (Yang et al., 2022, 2023; Wang et al., 2024; Zhu et al., 2023) employ contrastive learning to mitigate potential knowledge noise and identify informative knowledge connections.
@@ 2.2. LLMs for Recommendation
In light of the emergence of large language models and their remarkable achievements in the field of NLP, scholars have begun to explore the potential application of LLMs in recommender systems (Wu et al., 2023; Zhao et al., 2023b; Fan et al., 2023; Chen et al., 2023). Due to the powerful reasoning capabilities and extensive world knowledge of LLMs, they have been already naturally applied to zero-shot (Hou et al., 2024; He et al., 2023; Wang and Lim, 2023) and few-shot recommendation scenarios (Bao et al., 2023; Lin et al., 2024b). In these studies, LLMs are directly used as a recommendation model (Zhao et al., 2023a; Li et al., 2023), where the output of LLMs is expected to offer a reasonable recommendation result (Wu et al., 2024b). However, when the dataset is sufficiently large, their performance often falls short of that achieved by traditional recommendation models. Another line of research involves leveraging LLMs as feature extractors. These methods (Ren et al., 2024; Zhu et al., 2024; Hu et al., 2024; Wu et al., 2024a; Acharya et al., 2023; Kim et al., 2024; Wei et al., 2024; Ren et al., 2024) generate intermediate decision results or semantic embeddings of users and items, which are then input into traditional recommendation models to produce the final recommendations. Unlike existing methods, our approach aims to leverage the extensive knowledge and reasoning capabilities of LLMs to understand KGs and transform them into semantic embeddings, thereby addressing existing issues in KG-based recommender systems.
@@ 3. Preliminaries
  • Knowledge Graph. We capture real-world knowledge about items using a heterogeneous graph composed of triplets, represented as $\mathcal{G}_{k}=\{(h,r,t)\}$. In this context, $h$ and $t$ are knowledge entities belonging to the set $\mathcal{E}$, while $r$ is a relation from the set $\mathcal{R}$ that links them, as exemplified by the triplet (James Cameron, directed, Titanic). Notably, the item set is a subset of the entity set, denoted as $\mathcal{V}\subset\mathcal{E}$. This form of knowledge graph enables us to model the intricate relationships between items and entities.
@@ 4.1.2. Rerieval-based Global KG Utilization
This section introduces the utilization of global KG information. In recommendation scenarios, items that are distant in the KG can still have close semantic associations. However, the number of KG nodes increases exponentially with the number of hops, and the input length of LLMs is limited by the number of tokens. This makes it impractical to input the entire KG into an LLM directly. To address this challenge, we propose a method called retrieval-based global KG utilization.
@@
From a high-level perspective, we transform high-order associations between items in the KG into direct semantic connections on the constructed item-item graph $\mathcal{G}_{v}$. Based on this foundation, it is essential to retrieve items that are semantically strongly related to the given item, as items with lower semantic relevance may introduce noise. Specifically, given an item $v_{i}$, we rank all other items $v_{j}\in\mathcal{V}$ where $v_{j}\neq v_{i}$ in descending order based on the semantic similarity $r_{(v_{i},v_{j})}$. Subsequently, we retrieve the top-$k$ items with the highest similarity scores, forming the neighbor set of $v_{i}$: $\mathcal{N}_{k}(v_{i})$, where $0<k<|\mathcal{V}|$ is an adjustable hyperparameter, representing the number of retrieved neighbors.
In this manner, we filter out items with low semantic associations to the current item in the entire KG. Traditional KG-based recommendation methods aggregate items through layer-by-layer information propagation on the KG. High-order relations require many propagation and aggregation steps to be captured. In contrast, our retrieval-based method can directly recall strongly semantically related neighbors of any order across the entire KG. The retrieved neighbors are then leveraged to enhance item representations, which will be introduced in Section 4.3.2.
@@ 4.2. User Preference Comprehension
4.3. Retrieval-Augmented Representation
@@ 4.4. User-Item Modeling
where $\hat{y}_{uv}$ is the predicted probability of user $u$ interacting with item $v$, $\mathbf{h}_{u}$ is the representation for user $u$, $\mathbf{h}_{v}^{\prime}$ is the augmented representation for item $v$, and $\mathcal{F}$ denotes the function of the recommendation model.
@@ 4.5. Model Training
Our approach can be divided into two stages. In the first stage, we employ the LLM to comprehend the KGs, generating corresponding semantic embeddings for each item and user, denoted as $\mathbf{s}_{v}$ and $\mathbf{s}_{u}$, respectively. In the second stage, these semantic embeddings are integrated into the recommendation model to enhance its performance. Only the second stage necessitates supervised training, where we adopt the widely-used Bayesian Personalized Ranking (BPR) loss: